What plus -1 equals -20

Found 2 solutions by AnlytcPhil, ikleyn:
Answer by AnlytcPhil(1810)


What plus -1 equals -20?

Let the answer be x:

x + (-1) = -20

Add (+1) to both sides of the equation):

x + (-1) + (+1) = -20 + (+1)
              x = -20 + 1
              x = -19

The answer is -19.

Edwin

Answer by ikleyn(53751)

.

Informally, when you add -1 to a number, you shift the number one unit to the left on the number line.

So, if your shifted number is -20 after adding -1, it means that your original number was -19.

Thus, -19 is your answer in this problem.